this one is tough.. you are making a coffee table with a glass top surrounded by a cherry border. The glass is 3 feet by 3 feet. You want the cherry border to be uniform width. you have 7 square feet of cherry to make the border. set up an equation that would be used to solve this. what should the width of the border be?
if you could please help me.. it would be greatly appreciated. PLEASE! :]

Answer by edjones(7311)

A=3*3=9
P=4*3=12
(x+3)^2-9=7
x^2+6x+9-9=7
x^2+6x-7=0
(x+7)(x-1)=0
x=-7 not an answer,
x=1 ft Answer
Ed